# Unveiling the Pitfalls of Knowledge Editing for Large Language Models

### v3

@@ 2.2.2 Evaluation
$\mathrm{CS}=\mathbb{I}(p_{f_{\theta^{{}^{\prime}}}}(k_{n})>p_{f_{\theta^{{}^{\prime}}}}(k_{o})),$ (7)
@@
$\mathrm{Succ}=\mathbb{I}(p_{f_{\theta^{{}^{\prime}}}}(o\mid(s,r))>p_{f_{\theta}}(o\mid(s,r))).$ (9)
@@ 2.3.2 Evaluation
$\mathrm{IR}=\sum\limits_{\begin{subarray}{c}o\in\mathrm{Obj}\\ o\neq o_{1}\end{subarray}}\mathbb{I}(p_{f_{\theta}}(o\mid(s,r))>p_{f_{\theta^{{}^{\prime}}}}(o\mid(s,r))),$ (12)
@@
$\mathrm{FR}=\mathbb{I}(\mathrm{IR}>0.5).$ (13)
@@ A Simple Solution: Multi-Label Edit (MLE)
Input: James J. Lovelace was educated at _. (a) Easy
Input: James J. Lovelace was educated at _. (a) Easy
Input: Uri Avnery received award(s), such as _. (b) Hard
@@ A.1.2 RoundEdit
Easy
1-5 EducatedAt
James J. Lovelace was educated at _.
Cranbrook Educational Community
Target: United States Army Command and General Staff College, United States Military Academy, Indiana University, Salve Regina University, Naval War College
Hard
1-5 AwardReceived
Uri Avnery received award(s), such as _.
Hero of the Mongolian People’s Republic
Target: Sokolov Award, Right Livelihood Award, Blue Planet Award, Carl von Ossietzky Prize, Bruno Kreisky Award for Services to Human Rights
@@ Appendix C Details of Multi-Label Edit
To address the issue of Knowledge Distortion in knowledge editing, we introduce a novel approach termed Multi-Label Editing (MLE). This method diverges from conventional knowledge editing. Our objective is to simultaneously edit groups of knowledge that possess the same $(s,r)$ into the model, leveraging MEMIT’s multi-editing capabilities during the process. For instance, when we make an edit Joe Biden was born in California normal-$\rightarrow$ Florida, we can consider the differences in the knowledge structures of California and Florida, recall their related semantics within the model (such as America), and then regard ’Florida’ and ’America’ as labels utilizing MLE to edit.

### v4

@@ 2.2.2 Evaluation
$\mathrm{CS}=\mathbbm{1}\{p_{f_{\theta^{{}^{\prime}}}}(k_{n})>p_{f_{\theta^{{}^{\prime}}}}(k_{o})\},$ (7)
@@
$\mathrm{Succ}=\mathbbm{1}\{p_{f_{\theta^{{}^{\prime}}}}(o^{*}\mid(s,r))>p_{f_{\theta^{{}^{\prime}}}}(o\mid(s,r))\}.$ (9)
@@ 2.3.2 Evaluation
$\mathrm{IR}=\dfrac{1}{\left|\mathrm{Obj}\right|-1}\sum\limits_{\begin{subarray}{c}o\in\mathrm{Obj}\\ o\neq o_{1}\end{subarray}}\mathbbm{1}\{p_{f_{\theta}}(o\mid(s,r))>p_{f_{\theta^{{}^{\prime}}}}(o\mid(s,r))\},$ (12)
@@
$\mathrm{FR}=\mathbbm{1}\{\mathrm{IR}>0.5\}.$ (13)
@@ A Simple Solution: Multi-Label Edit (MLE)
Input: James J. Lovelace was educated at _.
(a) Easy
Input: Uri Avnery received award(s), such as _.
(b) Hard
@@ A.1.2 RoundEdit
Easy 1-5 EducatedAt James J. Lovelace was educated at _. Cranbrook Educational Community Target: United States Army Command and General Staff College, United States Military Academy, Indiana University, Salve Regina University, Naval War College
Hard 1-5 AwardReceived Uri Avnery received award(s), such as _. Hero of the Mongolian People’s Republic Target: Sokolov Award, Right Livelihood Award, Blue Planet Award, Carl von Ossietzky Prize, Bruno Kreisky Award for Services to Human Rights
@@ Appendix C Details of Multi-Label Edit
To address the issue of Knowledge Distortion in knowledge editing, we introduce a novel approach termed Multi-Label Editing (MLE). This method diverges from conventional knowledge editing. Our objective is to simultaneously edit groups of knowledge that possess the same $(s,r)$ into the model, leveraging MEMIT’s multi-editing capabilities during the process. For instance, when we make an edit Joe Biden was born in California $\rightarrow$ Florida, we can consider the differences in the knowledge structures of California and Florida, recall their related semantics within the model (such as America), and then regard ’Florida’ and ’America’ as labels utilizing MLE to edit.